# Discarding Lavrentiev’s Gap in Non-automonous and Non-Convex Variational Problems

### v1

@@ 8.2 The boundary approximation on a special open set
where the last inequality relies on the fact that $x_{N}\leq\psi(x^{\prime})+{\varkappa}_{n}$ and $y_{N}\geq{\varepsilon}_{n}/4$. This implies that $x-y$ belongs to the intersection of $\Lambda$ with the hypograph of $\psi$ where $u$ coincides with $\varphi$. This yields for every $t\in{\mathbb{R}}$,
@@
$v_{{\varepsilon}_{n},\delta_{n}}(x,t+L_{\varphi}|x-\widetilde{x}|)$ $={1}_{u}*_{x}{\varrho}_{{\varepsilon}_{n}}(x,t+L_{\varphi}|x-\widetilde{x}|)+\delta_{n}\alpha(t+L_{\varphi}|x-\widetilde{x}|)$
$={1}_{\varphi}*_{x}{\varrho}_{{\varepsilon}_{n}}(x,t+L_{\varphi}|x-\widetilde{x}|)+\delta_{n}\alpha(t+L_{\varphi}|x-\widetilde{x}|)\,.$
@@
and by (8.17), the left-hand side converges to 00. Since $({\overline{u}_{n}^{s}})_{n\in{\mathbb{N}}}$ converges to $u$ in $L^{1}(\Sigma)$, this implies that
@@
In the case $N=1$, a special open set is simply an interval $(a,b)$ and a neighbourhood of the boundary can be chosen as a (small) interval near the extremities $a$ or $b$. Let us focus for instance on the approximation near $a$. We replace the set $\Sigma$ by $(a,b)$. In analogy to the higher dimensional case considered above, this amounts to taking $\psi$ equal to the constant $a$ and $\Gamma=\{a\}$. We also replace ${\mathbb{R}}^{N-1}\times{\mathbb{R}}$ by ${\mathbb{R}}$. Accordingly, in Step 1, we remove all the $N-1$ dimensional balls. The rest of the proof is essentially the same, so we omit the details.
@@ 8.3 Gluing the local approximations
The same arguments leading to [Bousquet-Pisa, Lemma 6.3] imply that $D{\Phi_{n,\overline{\delta}}}=\lambda_{n}+h_{n,\overline{\delta}}{{\mathscr{H}}^{N+1}}$, where
@@
    Hence, one can apply [AmFuPa, Proposition 1.80] to the sequence of finite Borel measures on $U_{j}\cap\Omega$ defined by $\mu_{n}(A)=\int_{A}f(x,{\overline{u}_{n}^{{\varphi},j}},\nabla{\overline{u}_{n}^{{\varphi},j}}){\,{\rm d}x}$, to conclude that
@@
    where in the last equality we used that $\sum_{j=0}^{J}\Psi_{j}(x)=1$ and that $\Psi_{j}$ equals 00 outside $U_{j}$. On the other hand, using that ${\Phi_{n,\overline{\delta}}}$ converges to $\Phi_{{\infty},\overline{\delta}}$ in $L^{1}(\Omega\times{\mathbb{R}})$ and the fact that $\sup_{n}|D{\Phi_{n,\overline{\delta}}}|(\Omega\times{\mathbb{R}})<\infty$ by (8.23), we deduce that $D{\Phi_{n,\overline{\delta}}}$ weakly-* converges to $D\Phi_{{\infty},\overline{\delta}}$. The Reschetnyak lower semicontinuity theorem thus implies that
@@
Given any sequence $(\overline{\delta}_{i})_{i\in{\mathbb{N}}}$ tending to 00, we can thus find a subsequence $(n_{i})_{i\in{\mathbb{N}}}$ such that
@@
    We are now in position to prove (8.28) for we are in one of the following three cases. If $t+C|x-y|\geq M^{\prime}$, then by (8.22), ${\Phi_{n,\overline{\delta}}}(x,t+C|x-y|)=0\leq{\Phi_{n,\overline{\delta}}}(y,t)\,.$ If $t\leq-M^{\prime}$, then by (8.22) again, ${\Phi_{n,\overline{\delta}}}(x,t+C|x-y|)\leq 1+\overline{\delta}={\Phi_{n,\overline{\delta}}}(y,t)\,.$ Otherwise $-M^{\prime}\leq t\leq t+C|x-y|\leq M^{\prime}$ and in that case $\overline{\alpha}(t+C|x-y|)-\overline{\alpha}(t)=-\frac{C|x-y|}{2M^{\prime}}.$ In view of (8.30) and the definition of $C$, this yields (8.28), which implies that ${{u}_{i}^{s,{\varphi}}}$ is Lipschitz continuous on $\Omega$.
@@ 9.1 Proof of Theorem 3
We first extend $f$ on $\overline{\Omega}\times{\mathbb{R}}\times{\mathbb{R}}^{N}$ by setting $f(x,t,\xi)=0$ for every $x\in\partial\Omega$ and every $(t,\xi)\in{\mathbb{R}}\times{\mathbb{R}}^{N}$. Since $\Omega$ is bounded and Lipschitz, there exists an open set $\Omega_{0}\subset{\mathbb{R}}^{N}$ such that $\partial\Omega\subset\Omega_{0}$ and a biLipschitz map $\Pi_{0}:\Omega_{0}\to\Omega_{0}$ which agrees with the identity on $\partial\Omega$, and such that $\Pi_{0}(\Omega_{0}\setminus\overline{\Omega})=\Omega_{0}\cap\Omega$, see e.g. [Conti-Focardi-Iurlano, Theorem 3.1.]. We then set $\Lambda\coloneqq\Omega\cup\Omega_{0}$ and
@@
By the Chebyshev inequality |$|\{u\geq n\}|\leq{n^{-p^{*}}}\|u\|_{L^{p^{*}}(\Omega)}^{p^{*}},$ so the right-hand side is finite since $u\in W^{1,p}(\Omega)\subset L^{p^{*}}(\Omega)$. Hence,
@@
We first consider the case $p=1$. Let $f$ and $u$ as in the statement. By Lemma 9.2 and a diagonal argument, we may assume that $u\in L^{\infty}(\Omega)$. Using Lemma 9.1, we introduce an extension $f_{1}$ of $f$ on $\Lambda\times{\mathbb{R}}\times{\mathbb{R}}^{N}$, which satisfies ($H^{\textrm{conv}}$) on $\Lambda\times{\mathbb{R}}\times{\mathbb{R}}^{N}$. Note that by [Ekeland-Temam, Theorem VIII.1.3 and Lemma VIII.1.2] for $\nabla u\in L^{1}(\Lambda)$ there exists a superlinear convex function $h:{\mathbb{R}^{N}}\to{[0,\infty)}$ such that $h(\nabla u)\in L^{1}(\Lambda)$ and $h(0)=0$. Let $g_{0}$ and $g$ be the functions defined for a.e. $x\in\Lambda$ and all $(t,\xi)\in{\mathbb{R}}\times{\mathbb{R}^{N}}$ by
@@
The last equality above and the weak convergence of $(u_{n})_{n\in{\mathbb{N}}}$ in $W^{1,1}(\Omega)$ then imply the strong convergence in $W^{1,1}(\Omega)$, see e.g. [GMS-2, Section 1.3.4, Proposition 1]. Since $g_{0}$ is non-negative, it follows from the first equality that $(g_{0}(\cdot,u_{n},\nabla u_{n}))_{n\in{\mathbb{N}}}$ converges to $g_{0}(\cdot,u,\nabla u)$ in $L^{1}(\Omega)$. We note that
@@
9.2 Proofs of Theorems 2 and 4
Having proven Theorem 3, we may now infer Theorems 2 and 4.
@@ B Examples
Let us take any $\cal{k_{2}}>0$ and consider $x,t,\xi,{\varepsilon}$ such that ${(f_{\text{o}})}^{-}_{B(x,{\varepsilon})}(t,\xi)+|\xi|^{N}\leq{\cal{k}_{2}}{{\varepsilon}^{-N}}$. We get that $|\xi|\leq\min\left(\psi_{0}^{-1}({\cal{k}}_{2}{\varepsilon}^{-N}),{\cal{k}}_{2}^{1/N}{\varepsilon}^{-1}\right)$. Fix any $y\in B(x,{\varepsilon})$ and consider the function $C(\cdot)$ coming from the definition of ${\mathcal{Z}}_{\omega(\cdot,\cdot)}$. If $\frac{{\varepsilon}^{-N}}{\psi_{1}(\psi_{0}^{-1}({\cal{k}}_{2}{\varepsilon}^{-N}))}\leq\frac{\psi_{0}({\cal{k}}_{2}^{1/N}{\varepsilon}^{-1})}{\psi_{1}({\cal{k}}_{2}^{1/N}{\varepsilon}^{-1})}$, we use the inequality ${|\xi|\leq{\cal{k}}_{2}^{1/N}{\varepsilon}^{-1}}$ and condition (2.7) to get
$\frac{f_{\text{o}}(x,t,\xi)}{f_{\text{o}}(y,t,\xi)}\leq 1+C(t)+{C(t)}\omega({\varepsilon})\frac{\psi_{1}({\cal{k}}_{2}^{1/N}{\varepsilon}^{-1})}{\psi_{0}({\cal{k}}_{2}^{1/N}{\varepsilon}^{-1})}\leq 1+C(t)+{C(t)}\widetilde{C}(\cal{k}_{2})\,.$ (B.6)
On the other hand, if $\frac{{\varepsilon}^{-N}}{\psi_{1}(\psi_{0}^{-1}({\cal{k}}_{2}{\varepsilon}^{-N}))}\geq\frac{\psi_{0}({\cal{k}}_{2}^{1/N}{\varepsilon}^{-1})}{\psi_{1}({\cal{k}}_{2}^{1/N}{\varepsilon}^{-1})}$, we use the inequality $|\xi|\leq\psi_{0}^{-1}({\cal{k}}_{2}{\varepsilon}^{-N})$ and condition (2.7) to get
$\frac{f_{\text{o}}(x,t,\xi)}{f_{\text{o}}(y,t,\xi)}\leq 1+C(t)+{C(t)}\omega({\varepsilon})\frac{\psi_{1}(\psi_{0}^{-1}({\cal{k}}_{2}{\varepsilon}^{-N}))}{{\cal{k}}_{2}{\varepsilon}^{-N}}\leq 1+C(t)+{C(t)}\widetilde{C}(\cal{k}_{2})\,.$ (B.7)
@@
Example B.3 (Koch, Ruf, Schäffner [Lukas23]).
@@
The convexity of $f_{\text{W}}$ with respect to $\xi$ follows from the fact that the function $r\mapsto\exp(r^{q})$ is convex on $[r_{*},\infty)$ and its derivative at $r=r_{*}$ is equal to $\exp(r_{*}^{q})r_{*}^{-1}$. It is also clear that $f_{\text{W}}\leq f_{\text{D}}\leq f_{\text{W}}+\|a\|_{L^{\infty}}\exp(r^{q}_{*})$ and also that $f_{\text{W}}$ satisfies (2.1), as $f_{\text{W}}(x,t,0)\leq\|a\|_{L^{\infty}}$. We now aim at showing that $f_{\text{W}}$ satisfies ($H^{\textrm{iso}}$). By similar computations to those in Example B.2, we only have to prove that $\omega$ satisfies (2.7) with
@@
As ${\varkappa}>q\min(1,N/p)$, at least one of the expressions above tends to 00 as $s$ tends to 00, and therefore, (2.7) is satisfied for small $s$. For larger $s$, (2.7) is clear as the left-hand side is bounded from above while the right-hand side is bounded from below by positive constants independent of $s$. Hence, the function $f_{\text{W}}$ satisfies ($H^{\textrm{iso}}$). From Theorem 4 and Proposition A.1, we can now infer the absence of the Lavrentiev phenomenon for the functional $u\mapsto\int_{\Omega}f_{\text{D}}(x,u,\nabla u){\,{\rm d}x}$. ∎

### v3

@@ 8.2 The boundary approximation on a special open set
where the last inequality relies on the fact that $x_{N}\leq\psi(x^{\prime})+{\varkappa}_{n}$ and $y_{N}\geq{\varepsilon}_{n}/4$. This implies that xyx-y belongs to the intersection of $\Lambda$ with the hypograph of $\psi$ where $u$ coincides with $\varphi$. This yields for every $t\in{\mathbb{R}}$,
@@
[…]
@@
and by (8.17), the left-hand side converges to 0. Since $({\overline{u}_{n}^{s}})_{n\in{\mathbb{N}}}$ converges to $u$ in $L^{1}(\Sigma)$, this implies that
@@
In the case $N=1$, a special open set is simply an interval $(a,b)$ and a neighbourhood of the boundary can be chosen as a (small) interval near the extremities $a$ or $b$. Let us focus for instance on the approximation near $a$. We replace the set $\Sigma$ by $(a,b)$. In analogy to the higher dimensional case considered above, this amounts to taking $\psi$ equal to the constant $a$ and $\Gamma=\{a\}$. We also replace ${\mathbb{R}}^{N-1}\times{\mathbb{R}}$ by $\mathbb{R}$}. Accordingly, in Step 1, we remove all the N1N-1 dimensional balls. The rest of the proof is essentially the same, so we omit the details.
@@ 8.3 Gluing the local approximations
The same arguments leading to [16, Lemma 6.3] imply that $D{\Phi_{n,\overline{\delta}}}=\lambda_{n}+h_{n,\overline{\delta}}{{\mathscr{H}}^{N+1}}$, where
@@
    Hence, one can apply [2, Proposition 1.80] to the sequence of finite Borel measures on $U_{j}\cap\Omega$ defined by $\mu_{n}(A)=\int_{A}f(x,{\overline{u}_{n}^{{\varphi},j}},\nabla{\overline{u}_{n}^{{\varphi},j}}){\,{\rm d}x}$, to conclude that
@@
    where in the last equality we used that $\sum_{j=0}^{J}\Psi_{j}(x)=1$ and that $\Psi_{j}$ equals 0 outside $U_{j}$. On the other hand, using that $\Phi_{n,\overline{\delta}}$} converges to $\Phi_{{\infty},\overline{\delta}}$ in $L^{1}(\Omega\times{\mathbb{R}})$ and the fact that $\sup_{n}|D{\Phi_{n,\overline{\delta}}}|(\Omega\times{\mathbb{R}})<\infty$ by (8.23), we deduce that $D{\Phi_{n,\overline{\delta}}}$ weakly-* converges to $D\Phi_{{\infty},\overline{\delta}}$. The Reschetnyak lower semicontinuity theorem thus implies that
@@
Given any sequence $(\overline{\delta}_{i})_{i\in{\mathbb{N}}}$ tending to 0, we can thus find a subsequence $(n_{i})_{i\in{\mathbb{N}}}$ such that
@@
    We are now in position to prove (8.28) for we are in one of the following three cases. If $t+C|x-y|\geq M^{\prime}$, then by (8.22), ${\Phi_{n,\overline{\delta}}}(x,t+C|x-y|)=0\leq{\Phi_{n,\overline{\delta}}}(y,t)\,.$ If $t\leq-M^{\prime}$, then by (8.22) again, ${\Phi_{n,\overline{\delta}}}(x,t+C|x-y|)\leq 1+\overline{\delta}={\Phi_{n,\overline{\delta}}}(y,t)\,.$ Otherwise $M^{\prime}\leq t\leq t+C|x-y|\leq M^{\prime}$ and in that case $\overline{\alpha}(t+C|x-y|)-\overline{\alpha}(t)=-\frac{C|x-y|}{2M^{\prime}}.$ In view of (8.30) and the definition of $C$, this yields (8.28), which implies that ${u}_{i}^{s,{\varphi}}$} is Lipschitz continuous on $\Omega$.
@@ 9.1 Proof of Theorem 3
We first extend $f$ on $\overline{\Omega}\times{\mathbb{R}}\times{\mathbb{R}}^{N}$ by setting $f(x,t,\xi)=0$ for every $x\in\partial\Omega$ and every $(t,\xi)\in{\mathbb{R}}\times{\mathbb{R}}^{N}$. Since $\Omega$ is bounded and Lipschitz, there exists an open set $\Omega_{0}\subset{\mathbb{R}}^{N}$ such that $\partial\Omega\subset\Omega_{0}$ and a biLipschitz map $\Pi_{0}:\Omega_{0}\to\Omega_{0}$ which agrees with the identity on $\partial\Omega$, and such that $\Pi_{0}(\Omega_{0}\setminus\overline{\Omega})=\Omega_{0}\cap\Omega$, see e.g. [26, Theorem 3.1.]. We then set $\Lambda\coloneqq\Omega\cup\Omega_{0}$ and
@@
By the Chebyshev inequality $|\{u\geq n\}|\leq{n^{-p^{*}}}\|u\|_{L^{p^{*}}(\Omega)}^{p^{*}},$ so the right-hand side is finite since $u\in W^{1,p}(\Omega)\subset L^{p^{*}}(\Omega)$. Hence,
@@
The assumption ($H^{\textrm{conv}}$) and ($H^{\textrm{iso}}$) enjoy the following stability property.
@@
We first consider the case $p=1$. Let $f$ and $u$ as in the statement. By Lemma 9.2 and a diagonal argument, we may assume that $u\in L^{\infty}(\Omega)$. Using Lemma 9.1, we introduce an extension $f_{1}$ of $f$ on $\Lambda\times{\mathbb{R}}\times{\mathbb{R}}^{N}$, which satisfies ($H^{\textrm{conv}}$) on $\Lambda\times{\mathbb{R}}\times{\mathbb{R}}^{N}$. Note that by [31, Theorem VIII.1.3 and Lemma VIII.1.2] for $\nabla u\in L^{1}(\Lambda)$ there exists a superlinear convex function $h:{\mathbb{R}^{N}}\to{[0,\infty)}$ such that $h(\nabla u)\in L^{1}(\Lambda)$ and $h(0)=0$. Let $g_{0}$ and $g$ be the functions defined for a.e. $x\in\Lambda$ and all $(t,\xi)\in{\mathbb{R}}\times{\mathbb{R}^{N}}$ by
@@
The last equality above and the weak convergence of $(u_{n})_{n\in{\mathbb{N}}}$ in $W^{1,1}(\Omega)$ then imply the strong convergence in $W^{1,1}(\Omega)$, see e.g. [35, Section 1.3.4, Proposition 1]. Since $g_{0}$ is non-negative, it follows from the first equality that $(g_{0}(\cdot,u_{n},\nabla u_{n}))_{n\in{\mathbb{N}}}$ converges to $g_{0}(\cdot,u,\nabla u)$ in $L^{1}(\Omega)$. We note that
@@
9.2 Proofs of Theorems 2 and 5
Having proven Theorem 3, we may now infer Theorems 2 and 5.
@@ 9.2 Proofs of Theorems 2 and 5
It is enough to show that our regime implies ($H^{\textrm{conv}}$). Indeed, once we show ($H^{\textrm{conv}}$) the result holds due to Theorem 3.
@@ B Examples
We provide an example illustrating that condition ($H^{\textrm{conv}}$) and conditions ($H^{\textrm{iso}}$), ($H^{\rm iso}_{0}$) are essentially different when considered in the fully anisotropic setting.
@@
[…]
@@
Example B.3 (Koch, Ruf, Schäffner [43]).
@@
The convexity of $f_{\text{W}}$ with respect to $\xi$ follows from the fact that the function $r\mapsto\exp(r^{q})$ is convex on $r_{*},\infty)$ and its derivative at $r=r_{*}$ is equal to $\exp(r_{*}^{q})r_{*}^{-1}$. It is also clear that $f_{\text{W}}\leq f_{\text{D}}\leq f_{\text{W}}+\|a\|_{L^{\infty}}\exp(r^{q}_{*})$ and also that $f_{\text{W}}$ satisfies (1.8), as $f_{\text{W}}(x,t,0)\leq\|a\|_{L^{\infty}}$. We now aim at showing that $f_{\text{W}}$ satisfies ($H^{\textrm{iso}}$). By similar computations to those in Example B.2, we only have to prove that $\omega$ satisfies (2.4) with
@@
As ${\varkappa}>q\min(1,N/p)$, at least one of the expressions above tends to 0 as $s$ tends to 0, and therefore, (2.4) is satisfied for small $s$. For larger $s$, (2.4) is clear as the left-hand side is bounded from above while the right-hand side is bounded from below by positive constants independent of $s$. Hence, the function $f_{\text{W}}$ satisfies ($H^{\textrm{iso}}$). From Theorem 5 and Proposition A.1, we can now infer the absence of the Lavrentiev phenomenon for the functional $u\mapsto\int_{\Omega}f_{\text{D}}(x,u,\nabla u){\,{\rm d}x}$. ∎